# Generating Fock-state superposition from coherent state by quantum measurement

### v1

@@ III.2 Superposed Fock-state generation
One can find that for $l=1$ (generally for an odd $l$), the parity of $N$ determines the output state is $\left|\psi_{+}\right\rangle$ or $\left|\psi_{-}\right\rangle$. In another word, after a sufficient large number of measurements, the resonator state will oscillate back and forth between $\left|\psi_{+}\right\rangle$ and $\left|\psi_{-}\right\rangle$ with an increasing fidelity. That alternate pattern can be observed by the blue dashed line and the red dot-dashed line in Fig. 5(a), which is consistent with Eq. (24). One can therefore employ a combination $\mathcal{F}_{+}+\mathcal{F}_{-}$ to characterize the performance of our protocol. At the end of each round of evolution and measurement, $\mathcal{F}_{+}+\mathcal{F}_{-}=\mathcal{F}_{+}(\mathcal{F}_{-})$ for an even (odd) $N$. In Fig. 5(a), the fidelity $\mathcal{F}_{+}+\mathcal{F}_{-}$ of the superposed state $(\left|0\right\rangle\pm\left|5\right\rangle)/\sqrt{2}$ is plotted with various strategy about the arrangement of $\tau$. In the same spirit of the Fock-state generation in Sec. III.1, the hybrid strategy displays a dramatic advantage over the uniform strategy. For example, the fidelity of the target state $\left|\psi_{+}\right\rangle$ reaches $99\%$ with only $N=10$ measurements under the strategy $\mathcal{S}_{3}^{(5)}$. In contrast, it is about $71\%$ if $\tau$ remains as $\tau_{5}^{(g)}$ with the same number of measurements. Figure 5(b) demonstrates the success probability of the qubit on its ground state $\left|g\right\rangle$ as a function of $N$. Again, the success probability using the hybrid strategy achieves the saturation value with a faster rate than the uniform strategy.
@@
The difference between uniform and hybrid strategies in the superposed-state generation can be unambiguously described by the density matrix of the resonator. For either $|\psi_{+}\rangle$ in Figs. 6(a) and 6(b) or $|\psi_{-}\rangle$ in Figs. 6(c) and 6(d), which can be respectively prepared by $N=10$ and $N=11$ rounds of measurements, it is found that the hybrid strategy is superior to the uniform one.
@@
Figure 7: Performance of our protocol in generating various superposed Fock states $(\left|0\right\rangle\pm\left|n\right\rangle)/\sqrt{2}$ with the hybrid strategy $\mathcal{S}_{3}^{(5)}$. (a): Fidelity of combination $\mathcal{F}_{+}+\mathcal{F}_{-}$. (b) Success probability. The system parameters are the same as Fig. 2.
@@
Fidelities for more superposed Fock states are plotted in Fig. 7(a). Similar to the result of the Fock-state generation, more rounds of evolution and measurement are required for a larger $n$. It costs $N=6,8,14,$ and $20$ rounds of measurements to prepare the superposed states $(\left|0\right\rangle+\left|n\right\rangle)/\sqrt{2}$ of $n=2,4,6$, and $8$, respectively, with a fidelity over $99\%$. Note the population distribution of any coherent state (the initial state of the resonator) determines that it is hard to find in the same time sufficient occupations on both ground state $|0\rangle$ and Fock state $|n\rangle$ with a large $n$. In contrast, if one focuses on the Fock-state generation, it is then always possible to find a coherent state with a peak occupation on the desired $|n\rangle$. As one can expect from Eq. (29), the success probability of $(|0\rangle+|8\rangle)/\sqrt{2}$ in Fig. 7(b) is much lower than that of $|8\rangle$ in Fig. 4(b). In particular, the former is about $5\%$ and the latter is about $14\%$. As expected, the success probability for the superposed state also decreases with increasing $n$.
@@ IV Superposed Fock-state generation in double modes
Replacing the ancillary qubit with a $V$-type qutrit, our protocol for generating a single-resonator Fock state can be extended to a double-resonator system with frequencies $\omega_{a}\neq\omega_{b}$, as shown in Fig. 8. The ground state, the middle state, and the highest-level state of the ancillary system are denoted by $\left|g\right\rangle$, $\left|e\right\rangle$, and $\left|h\right\rangle$, respectively. And the resonator mode $a$ ($b$) is only coupled to the transition $\left|g\right\rangle\leftrightarrow\left|h\right\rangle$ ($\left|g\right\rangle\leftrightarrow\left|e\right\rangle$) transition with a coupling strength $g_{a}$ ($g_{b}$). $\delta_{h}=\omega_{h}-\omega_{a}$ and $\delta_{e}=\omega_{e}-\omega_{b}$ are the relevant detunings. In the rotating frame with respect to $U_{0}=\exp\{i[\omega_{a}(a^{\dagger}a+|h\rangle\langle h|)+\omega_{b}(b^{\dagger}b+|e\rangle\langle e|)]t\}$, the full Hamiltonian reads,
@@
Figure 8: Extended model diagram for our superposed Fock-state generation protocol, in which a $V$-type three-level system (qutrit) is coupled to two nondegenerate resonator modes. The projective measurements $M_{g}=|g\rangle\langle g|$ performing on the ancillary qutrit can steer the double modes towards a multi-excitation Bell state.
Our target is to generate a double-mode superposed Fock-state $\left|\psi_{\pm}\right\rangle=c_{00}\left|0\right\rangle_{a}\left|0\right\rangle_{b}\pm c_{mn}\left|m\right\rangle_{a}\left|n\right\rangle_{b}$ (briefly labelled as $c_{00}\left|00\right\rangle\pm c_{mn}\left|mn\right\rangle$ in the following), which is essentially a general Bell state of multiple excitations. The qutrit is initially at the ground state $\left|g\right\rangle$. And the two resonators $a$ and $b$ are prepared in coherent states $\left|\alpha\right\rangle$ and $\left|\beta\right\rangle$, respectively. To obtain the target state with desired amplitudes $c_{00}$ and $c_{mn}$, it is straightforward to find that $\alpha=(c_{mn}m!/c_{00})^{1/2m}$ and $\beta=(c_{mn}n!/c_{00})^{1/2n}$, in a similar way to generating a single-mode superposed state.
After $N$ rounds of evolution and measurement with a constant period $\tau$ and the measurement operator $M_{g}=|g\rangle\langle g|$, the state of resonators reads
@@
where $\Pi_{g}(\tau)=\left\langle g\right|e^{-iH^{\prime}t}\left|g\right\rangle$ denotes the effective time-evolution operator for the target resonators. $\Pi_{g}(\tau)$ is diagonal in the double-mode Fock basis $\{\left|kk^{\prime}\right\rangle\equiv\left|k\right\rangle_{a}\left|k^{\prime}\right\rangle_{b}\}$. And $P_{g}(N)$ is the success probability that the qutrit remains at its ground state after $N$ measurements. To find a compact expression for the measurement operator, we set $\delta_{g}=\delta_{e}=\delta$. Then we have
@@
$\lambda_{kk^{\prime}}(\tau)$ $=e^{-i\frac{\delta\tau}{2}}\left(\cos\Omega_{kk^{\prime}}\tau+i\frac{\delta}{2\Omega_{kk^{\prime}}}\sin\Omega_{kk^{\prime}}\tau\right),$ (34)
@@
Using Eqs. (32) and (33), the diagonal part of the density matrix of the two resonator modes $a$ and $b$ in the basis $\{\left|kk^{\prime}\right\rangle\}$ after $N$ rounds of equally spaced evolution and measurement is found to be
@@
where $p_{kk^{\prime}}\equiv\left\langle kk^{\prime}\right|\rho_{a}(0)\otimes\rho_{b}(0)\left|kk^{\prime}\right\rangle$ is their initial population and the success probability turns out to be
@@
The population-reduction factor on the double-mode Fock-states $\left|kk^{\prime}\right\rangle$ is determined by the modulus square of the coefficients $|\lambda_{kk^{\prime}}(\tau)|^{2}$. The matrix elements of off-diagonal part reads
@@
where $C_{kk^{\prime},mm^{\prime}}\equiv\left\langle kk^{\prime}\right|\rho_{a}(0)\otimes\rho_{b}(0)\left|mm^{\prime}\right\rangle$ represents the coherence of the initial state.
Under the resonant condition $\delta=0$, it turns out that $\lambda_{00}(\tau)=1$, irrespective to the evolution period $\tau$. In this case, $\left|g00\right\rangle$ is decoupled from all the other states of the composite system. From Eqs. (36) and (38), the fidelity of the target states $\left|\psi_{\pm}\right\rangle$ is found to be
@@
To generate $\left|\psi_{\pm}\right\rangle=c_{00}\left|00\right\rangle\pm c_{mn}\left|mn\right\rangle$, or equivalently to preserve both the population on $\left|mn\right\rangle$ and the coherence between $\left|00\right\rangle$ and $\left|mn\right\rangle$, the evolution period $\tau$ for each round of evolution and measurement can be chosen as $\tau=l\tau_{mn}=l\pi/\Omega_{mn}$ with $l\in\mathbb{N}^{+}$ to guarantee the condition $|\lambda_{mn}(\tau)|=1$. When $l$ is odd, the parity of $N$ will determine whether or not a $\pi$ phase presents in the coherence $\left\langle 00\right|\mathcal{C}[\rho_{a}(N\tau)]\left|mn\right\rangle$, i.e., the outcome state is $|\psi_{+}\rangle$ or $|\psi_{-}\rangle$. And similar to the preceding situations, the asymptotic success probability is equivalent to the overlap between the target states and the initial state of the resonator system. It is found to be
@@
Figure 9: 2D population-reduction factors $|\lambda_{kk^{\prime}}(\tau)|^{2N}$ with $N=5$ in the double-mode Fock-state basis $\{\left|kk^{\prime}\right\rangle\}$ under various coupling strengths and measurement intervals. In (a) and (c), $g_{a}=0.03\omega_{b}$ and $g_{b}=0.05\omega_{b}$; and in (b) and (d): $g_{a}=0.05\omega_{b}$ and $g_{b}=0.03\omega_{b}$. In (a) and (b), $\tau=\tau_{44}$; and in (c) and (d): $\tau=3\tau_{44}$.
Figure 10: (a) Fidelities of the target state $(\left|00\right\rangle\pm\left|44\right\rangle)/\sqrt{2}$ after $N$ rounds of evolution and measurement under the uniform strategy with fixed $\tau=\tau_{44}$, $g_{a}=0.05\omega_{b}$, and $g_{b}=0.03\omega_{b}$ or the hybrid strategy $\mathcal{S}_{3}^{(5,15)}$, in which the coupling strengths $g_{a}=0.05\omega_{b}$ and $g_{b}=0.03\omega_{b}$ are modified to be $g_{a}=0.03\omega_{b}$ and $g_{b}=0.05\omega_{b}$ after $L=15$ rounds of measurements. (b) Success probability of finding the qutrit in its ground state after $N$ rounds.
@@
Different from the single-mode Fock-state generation in Sec. III.1, we have an extra degree of freedom in modulating the efficiency for generation of a double-mode state, since the double-mode population reduction factor $|\lambda_{kk^{\prime}}(\tau)|^{2N}$ actually has a two-dimensional profile, as shown in Fig. 9. The target state is supposed to be, e.g., $|\psi_{\pm}\rangle=c_{00}\left|00\right\rangle\pm c_{44}\left|44\right\rangle$. Although many states other than the desired Fock states $\left|00\right\rangle$ and $\left|44\right\rangle$ also present with a unit or close-to-unit population-reduction factor. One can find in Fig. 9 that these unwanted states living in the dark-blue areas are distributed along the lines with a slope $-g_{b}/g_{a}$ passing through the point $(4j^{2},4j^{2})$, $j\in\mathbb{N}^{+}$. Note in Figs. 9(a) and 9(b), we have $g_{a}/g_{b}=3/5$ and $g_{a}/g_{b}=5/3$, respectively. The two profiles are dramatically different from each other while their crossing points are the desired Fock states $\left|00\right\rangle$ and $\left|44\right\rangle$. A solution to reduce the populations on the unwanted states is then to alter the ratio of the two coupling strengths in the whole sequence of evolution-and-measurement rounds.
@@
Similar to the single-mode case (see Fig. 2), a longer evolution period $\tau$, i.e., a larger $l$, results in a more concentrated population-protected regions in Fock space, indicating a higher generation efficiency. However, it also gives rise to more number of population protected regions. These two features can be identified in Figs. 9(c) and 9(d) by comparing to Figs. 9(a) and 9(b), respectively. Thus in regard of the 2D profile of population-reduction ratio, the hybrid strategy for the single-mode case is updated to have various ratio of the coupling strengths $g_{a}/g_{b}$ and various period $\tau$. It can be relabelled as $\mathcal{S}^{(q,L)}_{l}$, $L>q$. The coupling ratio $g_{a}/g_{b}$ is modified from $5/3$ to $3/5$ after $L$ rounds of measurements; and the round period is $\tau=\tau_{mn}$ for $1\leq N\leq q$ and $L+1\leq N\leq L+q$ and becomes $\tau=l\tau_{mn}$ for the rest rounds. The uniform strategy with invariant $\tau=\tau_{mn}$ and $g_{a}/g_{b}=5/3$ can be labelled as $S_{1}^{(\infty,\infty)}$.
Figure 11: Performance of our protocol in generating various multi-excitation Bell states $(\left|00\right\rangle+\left|nn\right\rangle)/\sqrt{2}$ under various hybrid strategy. (a): Combination fidelity $\mathcal{F}_{+}+\mathcal{F}_{-}$. (b) Success probability. The system parameters are the same as Fig. 10.
@@
In Fig. 10(a), we present the fidelity of the target state $(\left|00\right\rangle\pm\left|44\right\rangle)/\sqrt{2}$ as a function of the measurement number $N$ under either the uniform strategy or the hybrid strategy $\mathcal{S}^{(5,15)}_{3}$. $\mathcal{F}_{+}$ and $\mathcal{F}_{-}$ manifest the same alternate pattern for $N$ of different parity as in Fig. 5(a). With respect to the combination $\mathcal{F}_{+}+\mathcal{F}_{-}$, the hybrid strategy $\mathcal{S}^{(5,15)}_{3}$ demonstrates a significant advantage over the uniform strategy. With $30$ rounds of measurements, the former can prepare a general Bell state with a fidelity over $99\%$, while the latter attains only $25\%$ in fidelity. Figure 10(b) indicates the cost of the target-state preparation, in which the success probability is stabilized around about $3\%$ after about $25$ rounds of evolution and measurement.
@@
In Fig. 11, we show the performance of our protocol in generating Bell states $(\left|00\right\rangle+\left|nn\right\rangle)/\sqrt{2}$ with $n=1$, $3$, and $5$ under various hybrid strategy. Under the same strategy $S_{3}^{(5,8)}$, the fidelities of the target states with $n=1$ and $n=3$ reach $99\%$ after $N=9$ and $N=14$ rounds of measurements. And their success probabilities approach $28\%$ and $6\%$, respectively. It is evident that it is harder to prepare a general Bell state with a larger $n$. The comparison between different strategies on the same target state $(\left|00\right\rangle+\left|55\right\rangle)/\sqrt{2}$ indicates that it is nontrivial to select a proper strategy in state generation. During the first $L$ rounds with a constant ratio $g_{a}/g_{b}=5/3$, the unwanted yet protected populations should be sufficiently suppressed. Otherwise, the performance as a whole will become less optimized. That could explain why the final fidelity under $S_{3}^{(5,8)}$ is lower than that under $S_{3}^{(5,15)}$. In the latter case, one can prepare $(\left|00\right\rangle+\left|55\right\rangle)/\sqrt{2}$ by $N=33$ rounds of measurements with a fidelity over $99\%$.

### v2

@@ III.2 Superposed Fock-state generation
One can find that for $l=1$ (generally for an odd $l$), the parity of $N$ determining the output state is $|\psi_{+}\rangle$ or $|\psi_{-}\rangle$. In other word, after a sufficiently large number of measurements, the resonator state will oscillate between $|\psi_{+}\rangle$ and $|\psi_{-}\rangle$ with increasing fidelity. That alternate pattern can be observed from the blue dashed line and the red dash-dotted line in Fig. 7(a), which is consistent with Eq. (26). One can therefore employ the sum $\mathcal{F}_{+}+\mathcal{F}_{-}$ to characterize the performance of our protocol. At the end of each round of evolution and measurement, $\mathcal{F}_{+}+\mathcal{F}_{-}=\mathcal{F}_{+}(\mathcal{F}_{-})$ for an even (odd) $N$. In Fig. 7(a), the fidelity $\mathcal{F}_{+}+\mathcal{F}_{-}$ of the superposed state $(|0\rangle\pm|5\rangle)/\sqrt{2}$ is plotted for various strategies for an arrangement of $\tau$. In the same spirit as the Fock-state generation in Sec. III.1, the hybrid strategy displays a dramatic advantage over the uniform strategy. For example, the fidelity of the target state $|\psi_{+}\rangle$ reaches $99\%$ with only $N=10$ measurements under the strategy $\mathcal{S}_{3}^{(5)}$. In contrast, it is about $71\%$ if $\tau$ remains as $\tau_{5}^{(g)}$ with the same number of measurements. Figure 7(b) demonstrates the success probability of the qubit in its ground state $|g\rangle$ as a function of $N$. Again, the success probability using the hybrid strategy saturates at a faster rate than the uniform strategy.
@@
The difference between uniform and hybrid strategies in the superposed-state generation can be unambiguously described by the density matrix of the resonator. For either $|\psi_{+}\rangle$ in Figs. 8(a) and 8(b) or $|\psi_{-}\rangle$ in Figs. 8(c) and 8(d), which can be respectively prepared by $N=10$ and $11$ rounds of measurements, respectively, it confirms that the hybrid strategy is superior to the uniform one.
@@
Figure 9: Performance of our protocol in generating various superposed Fock states $(|0\rangle\pm|n\rangle)/\sqrt{2}$ with the hybrid strategy $\mathcal{S}_{3}^{(5)}$. (a) Sum of fidelity $\mathcal{F}_{+}+\mathcal{F}_{-}$. (b) Success probability. The system parameters are the same as in Fig. 3.
@@
Fidelities for more superposed Fock states are given in Fig. 9(a). Similar to the Fock-state generation in the closed-system scenario, more rounds of evolution and measurement are required for a larger $n$. It requires $N=6,~{}8,~{}14,$ and $20$ rounds of measurements to prepare the superposed states $(|0\rangle\pm|n\rangle)/\sqrt{2}$, with $n=2,~{}4,~{}6$, and $8$, respectively, with a fidelity over $99\%$. Note that the population distribution of any coherent state (the initial state of the resonator) determines that it is hard to find significant occupations in both the ground state $|0\rangle$ and Fock state $|n\rangle$ with a large $n$ at the same time. In contrast, if one focuses on the Fock-state generation, it is then always possible to find a coherent state with a peak occupation in the desired $|n\rangle$. As one can expect from Eq. (31), the success probability of $(|0\rangle\pm|8\rangle)/\sqrt{2}$ in Fig. 9(b) is much lower than that of $|8\rangle$ in Fig. 5(b). In particular, the former is about $5\%$ and the latter is about $14\%$. As expected, the success probability for the superposed state also decreases with increasing $n$.
@@ IV Superposed Fock-state generation in double modes
Replacing the ancillary qubit with a V-type qutrit, our protocol for generating a single-resonator Fock state can be extended to a double-resonator system with frequencies $\omega_{a}\neq\omega_{b}$, as shown in Fig. 10. The ground state, the middle state, and the highest-level state of the ancillary system are denoted by $|g\rangle$, $|e\rangle$, and $|h\rangle$, respectively. The resonator mode $a$ ($b$) is only coupled to the transition $|g\rangle\leftrightarrow|h\rangle$ ($|g\rangle\leftrightarrow|e\rangle$) with a coupling strength $g_{a}$ ($g_{b}$). In addition, $\delta_{h}=\omega_{h}-\omega_{a}$ and $\delta_{e}=\omega_{e}-\omega_{b}$ are the relevant detunings. In the rotating frame with respect to $U_{0}=\exp\{i[\omega_{a}(a^{\dagger}a+|h\rangle\langle h|)+\omega_{b}(b^{\dagger}b+|e\rangle\langle e|)]t\}$, the full Hamiltonian reads,
@@
Figure 10: Extended model diagram for our superposed Fock state generation protocol, in which a V-type three-level system (qutrit) is coupled to two nondegenerate resonator modes. The projective measurements $M_{g}=|g\rangle\langle g|$ performed on the ancillary qutrit can steer the double modes towards a multi excitation Bell state.
Our target is to generate a double-mode superposed Fock state $|\psi_{\pm}\rangle=c_{00}|0\rangle_{a}|0\rangle_{b}\pm c_{mn}|m\rangle_{a}|n\rangle_{b}$ (briefly labeled as $c_{00}|00\rangle\pm c_{mn}|mn\rangle$ in the following), which is essentially a general Bell state of multiple excitations. The qutrit is initially in the ground state $|g\rangle$. The two resonators $a$ and $b$ are prepared in coherent states $|\alpha\rangle$ and $|\beta\rangle$, respectively. To obtain the target state with desired amplitudes $c_{00}$ and $c_{mn}$, it is straightforward to find that $\alpha=(c_{mn}m!/c_{00})^{1/2m}$ and $\beta=(c_{mn}n!/c_{00})^{1/2n}$, in a similar way to generating a single-mode superposed state.
After $N$ rounds of evolution and measurement with a constant period $\tau$ and the measurement operator $M_{g}=|g\rangle\langle g|$, the state of the resonators reads
@@
where $\Pi_{g}(\tau)=\langle g|e^{-iH^{\prime}t}|g\rangle$ denotes the effective time-evolution operator for the target resonators; $\Pi_{g}(\tau)$ is diagonal in the double-mode Fock basis $\{|kk^{\prime}\rangle\equiv|k\rangle_{a}|k^{\prime}\rangle_{b}\}$. In addition, $P_{g}(N)$ is the success probability that the qutrit remains at its ground state after $N$ measurements. To find a compact expression for the measurement operator, we set $\delta_{g}=\delta_{e}=\delta$. Then we have Yan and Jing (2022); Puri (2001)
@@
$\lambda_{kk^{\prime}}(\tau)$ $=e^{-i(\delta\tau/2)}\left(\cos\Omega_{kk^{\prime}}\tau+i\frac{\delta}{2\Omega_{kk^{\prime}}}\sin\Omega_{kk^{\prime}}\tau\right),$ (36)
@@
Using Eqs. (34) and (35), the diagonal part of the density matrix of the two resonator modes $a$ and $b$ in the basis $\{|kk^{\prime}\rangle\}$ after $N$ rounds of equally spaced evolution and measurement is found to be
@@
where $p_{kk^{\prime}}\equiv\langle kk^{\prime}|\rho_{a}(0)\otimes\rho_{b}(0)|kk^{\prime}\rangle$ is their initial population and the success probability turns out to be
@@
The population-reduction factor of the double-mode Fock states $|kk^{\prime}\rangle$ is determined by the modulus square of the coefficients $|\lambda_{kk^{\prime}}(\tau)|^{2}$. The matrix elements of the off-diagonal part reads
@@
where $C_{kk^{\prime},mm^{\prime}}\equiv\langle kk^{\prime}|\rho_{a}(0)\otimes\rho_{b}(0)|mm^{\prime}\rangle$ represents the coherence of the initial state.
Under the resonant condition $\delta=0$, it turns out that $\lambda_{00}(\tau)=1$, irrespective of the evolution period $\tau$. In this case, $|g00\rangle$ is decoupled from all the other states of the composite system. From Eqs. (38) and (40), the fidelity of the target states $|\psi_{\pm}\rangle$ is found to be
@@
To generate $|\psi_{\pm}\rangle=c_{00}|00\rangle\pm c_{mn}|mn\rangle$, or equivalently to preserve both the population on $|mn\rangle$ and the coherence between $|00\rangle$ and $|mn\rangle$, the evolution period $\tau$ for each round of evolution and measurement can be chosen as $\tau=l\tau_{mn}=l\pi/\Omega_{mn}$ with $l\in\mathbb{N}^{+}$ to guarantee the condition $|\lambda_{mn}(\tau)|=1$. When $l$ is odd, the parity of $N$ will determine whether or not a $\pi$ phase presents in the coherence $\langle 00|\mathcal{C}[\rho_{a}(N\tau)]|mn\rangle$, i.e., the outcome state is $|\psi_{+}\rangle$ or $|\psi_{-}\rangle$. Further, similar to the preceding situations, the asymptotic success probability is equivalent to the overlap between the target states and the initial state of the resonator system. It is found to be
@@
Figure 11: Two-dimensional population-reduction factors $|\lambda_{kk^{\prime}}(\tau)|^{2N}$ with $N=5$ in the double-mode Fock-state basis $\{|kk^{\prime}\rangle\}$ for various coupling strengths and measurement intervals, with (a) and (c), $g_{a}=0.03\omega_{b}$ and $g_{b}=0.05\omega_{b}$ and (b) and (d) $g_{a}=0.05\omega_{b}$ and $g_{b}=0.03\omega_{b}$ and (a) and (b) $\tau=\tau_{44}$ and (c) and (d) $\tau=3\tau_{44}$.
@@
Different from the single-mode Fock-state generation in Sec. III.1, we have an extra degree of freedom in modulating the efficiency for generation of a double-mode state, since the double-mode population reduction factor $|\lambda_{kk^{\prime}}(\tau)|^{2N}$ actually has a two-dimensional profile, as shown in Fig. 11. The target state is assumed to be $|\psi_{\pm}\rangle=c_{00}|00\rangle\pm c_{44}|44\rangle$. However, many states other than the desired Fock states $|00\rangle$ and $|44\rangle$ also present with a unit or close-to-unit population-reduction factor. One can find in Fig. 11 that these unwanted states existing in the dark-blue areas are distributed along the lines with a slope $-g_{b}/g_{a}$ passing through the points $(4j^{2},4j^{2})$, $j\in\mathbb{N}^{+}$. Note that Figs. 11(a) and 11(b), we have $g_{a}/g_{b}=3/5$ and $5/3$, respectively. The two profiles are dramatically different from each other while their crossing points are the desired Fock states $|00\rangle$ and $|44\rangle$. A solution to reduce the populations of the unwanted states is then to alter the coupling-strength ratio in the whole sequence of evolution-and-measurement rounds.
Figure 12: (a) Fidelities of the target state $(|00\rangle\pm|44\rangle)/\sqrt{2}$ after $N$ rounds of evolution and measurement under the uniform strategy with fixed $\tau=\tau_{44}$, $g_{a}=0.05\omega_{b}$, and $g_{b}=0.03\omega_{b}$ or the hybrid strategy $\mathcal{S}_{3}^{(5,15)}$, in which the coupling strengths $g_{a}=0.05\omega_{b}$ and $g_{b}=0.03\omega_{b}$ are modified to be $g_{a}=0.03\omega_{b}$ and $g_{b}=0.05\omega_{b}$ after $L=15$ rounds of measurements. (b) Success probability of finding the qutrit in its ground state after $N$ rounds.
@@
Similar to the single-mode case (see Fig. 3), a longer evolution period $\tau$, i.e., a larger $l$, results in a more concentrated population-protection region in Fock space, indicating a higher generation efficiency. However, it also gives rise to more protected regions. These two features can be identified in Figs. 11(c) and 11(d) by comparing to Figs. 11(a) and 11(b), respectively. Thus, in regard to the two-dimensional profile of the population-reduction ratio, the hybrid strategy for the single-mode case is updated to have various ratios of the coupling strengths $g_{a}/g_{b}$ and various periods $\tau$. It can be relabeled as $\mathcal{S}^{(q,L)}_{l}$, $L>q$. In the following evaluation, the coupling ratio $g_{a}/g_{b}$ is modified from $5/3$ to $3/5$ after $L$ rounds of measurements; the round period is $\tau=\tau_{mn}$ for $1\leq N\leq q$ and $L+1\leq N\leq L+q$ and becomes $\tau=l\tau_{mn}$ for the remaining rounds. The uniform strategy with invariant $\tau=\tau_{mn}$ and $g_{a}/g_{b}=5/3$ can be labeled as $S_{1}^{(\infty,\infty)}$. Note that the strategies apply to arbitrary $g_{a}/g_{b}$.
@@
In Fig. 12(a) we present the fidelity of the target state $(|00\rangle\pm|44\rangle)/\sqrt{2}$ as a function of the measurement number $N$ under either uniform strategy or hybrid strategy $\mathcal{S}^{(5,15)}_{3}$. Here $\mathcal{F}_{+}$ and $\mathcal{F}_{-}$ manifest the same alternate pattern for $N$ of different parity as in Fig. 7(a). With respect to the fidelity sum $\mathcal{F}_{+}+\mathcal{F}_{-}$, the hybrid strategy $\mathcal{S}^{(5,15)}_{3}$ demonstrates a significant advantage over the uniform strategy. With $30$ rounds of measurements, the former can prepare the Bell-like state with a fidelity over $99\%$, while the latter attains only $25\%$ in fidelity. Figure 12(b) indicates the cost of the target-state preparation, in which the success probability is stabilized around about $3\%$ after about $25$ rounds of evolution and measurement.
Figure 13: Performance of our protocol in generating various Bell-like states $(|00\rangle\pm|nn\rangle)/\sqrt{2}$ under various hybrid strategies. (a) Fidelity sum $\mathcal{F}_{+}+\mathcal{F}_{-}$. (b) Success probability. The system parameters are the same as in Fig. 12.
@@
In Fig. 13, we show the performance of our protocol in generating Bell-like states $(|00\rangle\pm|nn\rangle)/\sqrt{2}$ with $n=1$, $3$, and $5$ under various hybrid strategies. Under the same strategy $S_{3}^{(5,8)}$, the fidelities of the target states with $n=1$ and $3$ reach $99\%$ after $N=9$ and $16$ rounds of measurements. Their success probabilities approach $27\%$ and $5\%$, respectively. It is evident that it is harder to prepare a Bell-like state with a larger $n$. The comparison between different strategies for the same target state $(|00\rangle\pm|55\rangle)/\sqrt{2}$ indicates that it is nontrivial to select a proper strategy in state generation. During the first $L$ rounds with a constant ratio $g_{a}/g_{b}$, the unwanted yet protected populations should be sufficiently suppressed. Otherwise, the performance as a whole will become less optimized. That could explain why the final fidelity under $S_{3}^{(5,8)}$ is lower than that under $S_{3}^{(5,15)}$. In the latter case, one can prepare $(|00\rangle\pm|55\rangle)/\sqrt{2}$ in $N=33$ rounds of measurements with a fidelity over $99\%$.
Figure 14: Performance of our protocol for generating $(|00\rangle\pm|nn\rangle)/\sqrt{2}$ after $N$ rounds of evolution and measurement with various decoherence rate under the hybrid strategy $\mathcal{S}_{3}^{(5,15)}$ for $g_{a}=0.05\omega_{e}$, $g_{b}=0.03\omega_{e}$, and (a) and (c) $n=3$ and (b) and (d) $n=5$. (a) and (b): Fidelity of the target Bell-like states. (c) and (d): Success probability.
By taking all the relevant dissipation and dephasing channels into account, the Lindblad master equation for the two-resonator-qutrit system can be written as
$\dot{\rho}(t)=$ $-i[H^{\prime},\rho{(t)}]+\gamma_{ge}\mathcal{L}[|g\rangle\langle e|]\rho(t)+\gamma_{eh}\mathcal{L}[|e\rangle\langle h|]\rho(t)$ (43)
$+\gamma_{gh}\mathcal{L}[|g\rangle\langle h|]\rho(t)+\kappa_{a}\mathcal{L}[a]\rho(t)+\kappa_{b}\mathcal{L}[b]\rho(t)$
$+\gamma_{e,\phi}\mathcal{L}[|e\rangle\langle e|]\rho(t)+\gamma_{h,\phi}\mathcal{L}[|h\rangle\langle h|]\rho(t),$
where $\kappa_{i}$ with $i=a,b$ represents the decay rate of resonator mode $i$, $\gamma_{ij}$ with $i,j=g,e,h$ represents the dissipation rate from the qutrit level $|j\rangle$ to $|i\rangle$, and $\gamma_{i,\phi}$ with $i=e,h$ represents the dephasing rate of the qutrit level $|i\rangle$. Here we set $\kappa=\kappa_{a}=\kappa_{b}$, $\gamma=\gamma_{ge}=\gamma_{eh}=\gamma_{gh}=\gamma_{e,\phi}=\gamma_{h,\phi}$, and $\kappa=0.5\gamma$.
In Figs. 14(a) and 14(b) we present the fidelities of a two-mode Bell-like state $(|00\rangle\pm|nn\rangle)/\sqrt{2}$ with $n=3$ and $5$, respectively, in the presence of decoherence. Here the fidelity $\mathcal{F}=\mathcal{F}_{+}$ when $N$ is even and $\mathcal{F}=\mathcal{F}_{-}$ when $N$ is odd. It is found that the Bell-like state can be approximately obtained only for a low decoherence rate $\gamma/g_{a}=10^{-3}$. On performing more and more rounds of measurements, an optimized $N$ turns out to achieve the maximum fidelity. With the same $\gamma/g_{a}=10^{-2.5}$, the fidelity for generating $(|00\rangle\pm|33\rangle)/\sqrt{2}$ attains $\mathcal{F}\approx 94\%$ when $N=21$ and that for $(|00\rangle\pm|55\rangle)/\sqrt{2}$ attains $\mathcal{F}\approx 86\%$ when $N=29$. Over the peak value, the fidelity decreases with $N$. The success probabilities for $n=3$ and $5$, are shown in Figs. 14(c) and 14(d), respectively. A larger $n$ gives rise to a lower final success probability.